The first terms of an A.P and G.P are each 2/3. Their common and difference and

common ratio are each X and the sum of their first three terms are also equal.

Find the two possible values of X


1
Expert's answer
2021-10-04T16:17:34-0400
"a=b=\dfrac{2}{3}"

"a+a+x+a+2x=b+bx+bx^2"

"3(\dfrac{2}{3})+3x=\dfrac{2}{3}(1+x+x^2)"

"6+9x=2+2x+2x^2"

"2x^2-7x-4=0"

"D=(-7)^2-4(2)(-4)=81"

"x=\dfrac{7\pm \sqrt{81}}{2(2)}"

"x_1=\dfrac{7- 9}{4}=-\dfrac{1}{2}"

"x_2=\dfrac{7+ 9}{4}=4"


"x\in\{-\dfrac{1}{2}, 4\}"